Is 8/9 the multiplication inverse of 1 whole 1/8 ? Why or why not? 

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

1 whole 1/8=-7/8
[Multiplicative inverse⟹product should be 1]
According to the question,
8/9 × (-7/8) =-7/9 ≠ 1
Therefore, 8/9 is not the multiplicative inverse of1 whole 1/8

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We know that the product of the number and its multiplicative inverse is 1.
 1 whole 1/8 = 9/8
9/8 × 8/9 = 1
 
Therefore, 8/9 is the multiplicative inverse of 1 whole1/8.
 
Hope it helps!
Thank you ✌️